sketch the normal curve and find the area in each of the following cases 

1. between z=-2.76 and z= -1.2

2. To the right of z=1.31

3. To the left of z=0.35

4.to the right of z= -0.95


1
Expert's answer
2022-03-17T03:36:46-0400

The sought area is painted black

1

"P(-2.76<Z<-1.2)=P(Z<-1.2)-P(Z<-2.76)=0.11507-0.00289=0.11218"


2



"P(Z>1.31)=1-P(Z<1.31)=1-0.90490=0.0951"





"P(Z<0.35)=0.63683"



"P(Z>-0.95)=1-P(Z<-0.95)=1-0.17106=0.82894"